Suppose a credit card charges a 10% annual interest rate and recognizes the outstanding balance at the end of every month. How much would you end up paying for a 10,000 USD outstanding balance for three months? (Round the the nearest decimal) If the answer is A,B, C, or D, simply type the respective letter in the answer box. If the answer is E, write down the correct answer.

A. 200.8

B. 167.4

C. 250

D. 252.1

E. None of the above:
Answer #1

Solution

Answer-$252.1

Future value=Initial amount*(1+r)^n

where

r=rate of intrest per period=10/12=0.833333% per month

n=number of periods=3

Initial amount=10000

Future value=10000*(1+.00833333)^3

=10252.09

Thus the intrest he will end up paying=10252.09-10000=252.09

Answer=$252.1

If you are satisfied with the answer,please give a thumbs up